# Determining Molecular Ground State with Quantum Imaginary Time Evolution using Broken-Symmetry Wave Function

### v1

@@ Abstract
Index Terms:
Broken Symmetry Wave Function, Quantum Imaginary Time Evolution, Spin Operator, Open Shell System, Ground State Energy Estimation
@@ II-A Imaginary-time Schrödinger equation
$-\partial_{\tau}|\Phi(\tau)\rangle=\hat{H}|\Phi(\tau)\rangle.$ (1)
@@ II-B Spin operators
Spin is an intrinsic angular momentum without a classical analog, and it is described by a spin operator $\hat{\mathbf{S}}=(\hat{S}_{x},\hat{S}_{y},\hat{S}_{z})$. The spin operator components obey the usual angular momentum commutation relations $[\hat{S}_{j},\hat{S_{k}}]=i\epsilon_{jkl}\hat{S}_{l}$, where $\epsilon_{jkl}$ is the Levi-Civita symbol, $j,k,l=x,y,z$, and we work in units where $\hbar=1$. The spin magnitude is described by the square of the spin operator $\hat{\mathbf{S}}^{2}$, which commutes with components of the spin operator $[\hat{\mathbf{S}}^{2},\hat{S}_{i}]=0$, for $i=x,y,z$. This leads to a common set of eigenvectors for $\hat{\mathbf{S}}^{2}$ and one of the spin components, usually taken to be $\hat{S}_{z}$. These eigenvectors $|s,m\rangle$ are labeled by the spin $s$ and its projection along the z-direction $m$, and satisfy the following eigenvalue equations,
@@ V-A N2 using ITE by matrix exponentiation
Here, $n_{\mathrm{LUNO}+i}$ is the occupation number of the $i$-th lowest unoccupied natural orbital (LUNO$+i$). For convenience, we denote the diradical character $y^{\text{PUHF}}_{i}$ as $y$ throughout the rest of the paper. It ranges from 00 to $1$, a molecule with $y>0$ is considered to have open shell character which becomes more prominent as $y$ increases. At the extremes, $y=0$ for a closed shell molecule, and $y=1$ for a pure diradical.
@@ A-A Branching diagram functions
The branching diagram shown in Fig. 7 is a blueprint for constructing spin eigenfunctions [34]. It consists of circles on the vertices and diagonal upward ($45^{\circ}$) and downward ($-45^{\circ}$) lines connecting vertices. The upward (downward) line represents adding an electron spin to increase (decrease) spin quantum number. The number in the circle represents the number of possible pathways to arrive the vertex and it corresponds to the number of spin eigenfunctions for given $N$ and $S$. In this work, we use the branching diagram to calculate the expansion coefficients of the spin eigenfunctions for the BS wave function. In fact, the BS wave function of an $N$-spin system with magnetic quantum number $M$ can be written as the linear combinations of the spin eigenfunctions as follows:

### v2

@@ II-A Imaginary-time Schrödinger equation
$\partial_{\tau}|\Phi(\tau)\rangle=\hat{H}|\Phi(\tau)\rangle.$ (1)
@@ II-B Spin operators
Spin is an intrinsic angular momentum without a classical analog, and it is described by a spin operator $\hat{\mathbf{S}}=(\hat{S}_{x},\hat{S}_{y},\hat{S}_{z})$. The spin operator components obey the usual angular momentum commutation relations $\hat{S}_{j},\hat{S_{k}}]=i\epsilon_{jkl}\hat{S}_{l}$, where $\epsilon_{jkl}$ is the Levi-Civita symbol, $j,k,l=x,y,z$, and we work in units where $\hbar=1$. The spin magnitude is described by the square of the spin operator $\hat{\mathbf{S}}^{2}$, which commutes with components of the spin operator $\hat{\mathbf{S}}^{2},\hat{S}_{i}]=0$, for $i=x,y,z$. This leads to a common set of eigenvectors for $\hat{\mathbf{S}}^{2}$ and one of the spin components, usually taken to be $\hat{S}_{z}$. These eigenvectors $|s,m\rangle$ are labeled by the spin $s$ and its projection along the z-direction $m$, and satisfy the following eigenvalue equations,
@@ V-A N2 using ITE by matrix exponentiation
Here, $n_{\mathrm{LUNO}+i}$ is the occupation number of the $i$-th lowest unoccupied natural orbital (LUNO+i+i). For convenience, we denote the diradical character $y^{\text{PUHF}}_{i}$ as $y$ throughout the rest of the paper. It ranges from 0 to $1$, a molecule with $y>0$ is considered to have open shell character which becomes more prominent as $y$ increases. At the extremes, $y=0$ for a closed shell molecule, and $y=1$ for a pure diradical.
@@ A-A Branching diagram functions
The branching diagram shown in Fig. 7 is a blueprint for constructing spin eigenfunctions [34]. It consists of circles on the vertices and diagonal upward ($45^{\circ}$) and downward ($45^{\circ}$) lines connecting vertices. The upward (downward) line represents adding an electron spin to increase (decrease) spin quantum number. The number in the circle represents the number of possible pathways to arrive the vertex and it corresponds to the number of spin eigenfunctions for given $N$ and $S$. In this work, we use the branching diagram to calculate the expansion coefficients of the spin eigenfunctions for the BS wave function. In fact, the BS wave function of an $N$-spin system with magnetic quantum number $M$ can be written as the linear combinations of the spin eigenfunctions as follows: